# Conghui He4, Xiaoyang Wang6, Fan Wu1

### v1

@@ 6.2 Semi-Structured Data
Diverse Markup Language Formats. Markup languages encompass a wide spectrum, including HTML, XML, JSON, Markdown, LaTeX, and domain-specific standards such as Journal Article Tag Suite (JATS). However, most existing research remains concentrated on a few widely used formats, primarily including HTML, XML, and JSON which often derived from web-scraped data. This leaves domain-specific or less-studied markup languages, as well as proprietary formats in industry, underexplored.
Addressing this gap requires the creation of specialized benchmarks for languages such as JATS, LaTeX, and Markdown to evaluate tasks like structure-aware question answering and format conversion. Moreover, model architectures must evolve towards structure-aware tokenization and training paradigms that explicitly leverage the hierarchical organization of markup, rather than treating markup as plain text sequences.
Understanding Complex Structured Table Semi-structured tables often exhibit irregular relationships and intricate nested structures, which remain difficult for current LLMs to fully capture. While recent approaches attempt to encode table structures and leverage the semantic reasoning capacity of LLMs, these methods are usually tailored to specific table types (e.g., financial tables). It is quite challenging to handle diverse semi-structured tables with complex structures.
To address this, researchers have explored several directions. One line of work focuses on Structure-Aware Encodings, where the table’s hierarchical or graph-like organization (e.g., DOM trees for HTML tables or JSON schemas) is explicitly modeled and injected into LLMs. Another approach emphasizes Intermediate Representations, such as transforming semi-structured tables into canonical forms or domain-specific languages (DSLs) that better align with LLMs’ sequential reasoning ability. In addition, Retrieval-Augmented Methods leverage external knowledge bases or schema repositories to supplement missing semantics, enabling LLMs to ground their reasoning in richer contexts.
@@ 6.3 Unstructured Data
High-Level Chart Understanding. While significant progress has been made in basic chart recognition and data extraction, critical higher-order tasks such as chart-based fact checking, detection of visual or statistical misrepresentations (e.g., truncated axes, inconsistent scales, or fabricated trends), and cross-modal chart data verification remain largely unaddressed. These tasks demand not only accurate perception of visual elements but also reasoning about data semantics, statistical validity, and alignment with external knowledge or textual assertions. The absence of standardized benchmarks and evaluation protocols for such tasks further hinders systematic progress.
We advocate for the development of comprehensive frameworks that support critical chart literacy—enabling systems to not just “read” charts, but to audit them. This includes creating datasets with annotated errors, factual inconsistencies, and deceptive visual practices, as well as designing models capable of joint visual-semantic reasoning to validate the integrity and truthfulness of the graph.
@@
Inadequate Multimodal Reasoning for Diverse Chart Types. Charts are inherently multimodal artifacts that fuse structured visual encodings (e.g., glyphs, spatial arrangements, color mappings) with natural language (titles, axis labels, captions). Current models often treat chart understanding as a vision-only or simplified vision-language problem, failing to capture the nuanced interplay between graphical conventions, domain-specific semantics, and contextual information. This limitation becomes especially pronounced when handling complex or non-standard chart types (e.g., small multiples, Sankey diagrams, or interactive dashboards).
A holistic, end-to-end chart understanding pipeline that unifies low-level perception (e.g., element detection), mid-level parsing (e.g., data-value mapping), and high-level inference (e.g., trend interpretation or anomaly detection) is needed. Such a pipeline should be modular, chart-type-agnostic, and grounded in both visualization theory and real-world usage patterns, enabling robust performance across diverse downstream applications.
Computational Efficiency of Video Analysis. Video is long-form sequential data; fixed sampling or full-frame processing consumes large compute and wastes capacity. Compressing frames/tokens often sacrifices fine-grained temporal cues, hurting localization and detail recognition.
@@
To address this, we can develop representations that preserve fine-grained visual detail while offering controllable context length (e.g., temporal separator tokens, dynamic sampling, learnable token merging, temporal positional encodings). Integrate temporal modeling and token-aggregation objectives into training to achieve a better trade-off between efficiency and fidelity.
@@
Video Temporal Localization with Semantic Coherence. Precise temporal localization requires detecting sparse, transient cues (e.g., action start/end, sudden events) and maintaining semantic consistency across long sequences. Current methods still struggle with event-boundary detection, timestamp reasoning, and preventing erroneous cross-time associations.
To address this,we can improve fine-grained event detection and semantic continuity by incorporating temporally aware tokens and supervision objectives directly into LLM-friendly representations.
Video Multimodal Fusion and Scalability. Effectively time-aligning and fusing visual, audio, and textual streams is both modeling and engineering challenging; high-quality, fine-grained human annotation is costly and hard to scale. Domain-specific tasks (emotion, gesture, anomaly detection, etc.) often rely on specialized labels and bespoke pipelines, which limits generalization.
@@
To address this,we can combine modular architectures (lightweight temporal adapters, instruction-aware selectors), multi-agent data curation/reviewer pipelines, and large-scale synthetic video–text generation to improve annotation efficiency, domain adaptation, and instruction-conditioned reasoning. Encourage task-centered cooperative agents that decompose complex video-understanding problems into simpler subtasks.
@@
Generalizing Multimodal Understanding for Documents. While models are moving beyond OCR, a core challenge is generalizing deep multimodal understanding—the holistic integration of text, layout, and visuals—across diverse domains and languages. General-purpose models often struggle with the unique structural grammars of specialized documents, such as the dense layouts of scientific papers or non-linear paths in historical manuscripts. Similarly, true multilingual generalization is hampered by language-specific conventions, like right-to-left scripts, revealing a gap between claimed capabilities and robust real-world applicability.
A dual approach is needed. First, developing efficient, document-native architectures (e.g., DocLLM) that prioritize layout information over expensive visual encoding. Second, enhancing model versatility through layout-aware instruction tuning that teaches reasoning about structure at multiple granularities. This must be supported by creating large-scale, high-quality datasets covering a broader range of languages and professional domains to enable true generalization.
@@
Efficiency in Large-Scale Document Processing. Processing large volumes of long-form documents, such as legal archives, presents a trade-off between computational efficiency and analytical depth. The limited context windows of most LLMs force chunking strategies that can sever long-range dependencies, while preserving fine-grained spatial details across thousands of pages remains a major technical hurdle.
Future work should focus on adaptive context management and hierarchical document representations. Combining these with progressive reasoning strategies, where a model first builds a coarse understanding before focusing on details, can create a better balance between computational cost and high-fidelity analytical accuracy.
@@
Accuracy and Robustness in Program Vulnerability Detection. While LLMs show promise in software vulnerability detection, the accuracy of state-of-the-art methods remains unsatisfactory, with reported success rates around 67.6% for detection and only 20% for automated repair. The number of vulnerabilities continues to grow, with over 25,000 new CVEs reported in 2024 alone. Moreover, the performance of deep learning-based methods has been shown to degrade significantly on rigorously validated datasets and in the face of simple code modifications.
The development of novel frameworks that integrate Code Property Graphs (CPGs) with LLMs presents a promising path. Techniques like CPG-guided code slicing can reduce code size while preserving vulnerability-relevant context, leading to significant improvements in F1 scores for inter-procedural vulnerability detection.
@@
Program Analysis Complexity and Scope Limitations. The inherent complexity of vulnerabilities poses a significant challenge, with inter-procedural flaws being substantially harder to detect than intra-procedural ones. LLMs often struggle with uncommon CWE types, and their performance degrades when vulnerabilities span multiple code units. Most current LLM-based solutions are limited to the function level, a narrow scope that questions their effectiveness for class-level or repository-level analysis.
Future research should focus on developing LLM architectures that support multi-level code analysis, scaling from the function level to project and enterprise levels. Integrating Retrieval-Augmented Generation (RAG) techniques is crucial for effectively handling the vast contextual information present in large-scale codebases.
Reliability in Code Generation and Repair. Although LLMs offer advantages in generality and interpretability—providing intermediate reasoning for their findings—they still face significant challenges in terms of trust and accuracy, particularly for automated code repair.
@@
Establishing mechanisms for trust and collaboration with developers is essential. LLMs should be designed to signal uncertainty rather than providing potentially inaccurate outputs. Furthermore, developing multi-agent collaborative frameworks, with specialized agents for code analysis, vulnerability detection, and repair, could lead to more robust and reliable systems.
High Quality 3D Modality Representation. 3D data come as point clouds, meshes, voxels, implicit fields, and NeRFs, each with different sampling, topology, and attributes. Converting these formats into compact LLM-ready inputs forces trade-offs: either aggressive simplification that loses geometric detail, or token explosion that breaks context budgets. These inconsistencies cause downstream reasoning to see uneven geometry and appearance cues, complicating generalization across tasks and formats.
@@
To address this,we can develop compact, trainable encoders that convert raw 3D inputs into a small number of informative tokens suited for LLMs. Examples include vector-quantized mesh tokens, learned point-feature aggregators, and geometry-aware positional encodings. Jointly train these encoders with lightweight adapters using multiview-consistency, reprojection, and geometric-accuracy losses so tokenization preserves topology and metric cues while respecting LLM context budgets.
@@
Accurate 3D Geometric Grounding. LLMs lack native metric reasoning and do not handle coordinate frames, units, or precise spatial numerics reliably. The scarcity of richly labeled 3D corpora with exact numeric annotations hinders supervised learning for distances, poses, and geometric constraints. As a result, systems may give plausible qualitative descriptions but fail on quantitative queries, undermining performance in navigation, manipulation, simulation, and safety-critical use cases.
To address this,we can use domain adapters that encode coordinate conventions, units, and regression heads for distances/poses, trained with metric supervision from simulation/renderers and differentiable-rendering or physical-constraint losses; evaluate on numerically grounded benchmarks to ensure quantitative fidelity.
@@
Pluggable Modality Extention Ability. A critical challenge for heterogeneous data agents is supporting the ability to incorporate new and diverse data types without redesigning or retraining the entire system. Traditional architectures are often rigid, requiring substantial modifications to handle emerging modalities such as audio, video, or sensor data. Without flexible integration mechanisms, agents risk becoming obsolete when encountering new data formats, limiting their long-term applicability and scalability.
Future work should focus on designing modular and extensible frameworks that allow seamless addition of new modalities through independent encoders or adapters. Research directions include developing standardized interfaces for modality integration, adaptive weighting strategies to balance multiple modalities, and self-supervised alignment methods to ensure that newly added modalities are compatible with existing representation spaces. Such approaches will enable heterogeneous data agents to remain scalable, maintainable, and robust in the face of evolving and increasingly complex datasets.
Heterogeneous High-Level Data Analysis. A key challenge in heterogeneous data agents is performing high-level analysis that goes beyond simple retrieval or extraction to include reasoning, summarization, and complex cross-modal inference. Unlike low-level tasks, high-level analysis requires understanding interrelationships among multiple modalities, capturing context, and generating actionable insights or visualizations. Existing systems, such as XMODE nooralahzadeh2024explainable and wang2024interactive , demonstrate partial solutions through step-wise query decomposition or adaptive modality weighting, yet they still struggle with disorganized or large-scale heterogeneous datasets, limiting their ability to provide fully automated, context-aware analyses.
Future research should focus on designing agents capable of robust, end-to-end high-level reasoning across heterogeneous modalities. Potential directions include leveraging multimodal embeddings for joint inference, and combining symbolic reasoning with neural models to ensure interpretability and correctness. Additionally, adaptive task planning mechanisms could enable agents to dynamically decompose complex queries into subtasks and synthesize results across modalities, thereby enhancing their capability for comprehensive, high-level data analysis in real-world scenarios.

### v2

@@ 6.2 Semi-Structured Data
Diverse Markup Language Formats. Markup languages encompass a wide spectrum, including HTML, XML, JSON, Markdown, LaTeX, and domain-specific standards such as Journal Article Tag Suite (JATS). However, most existing research remains concentrated on a few widely used formats, primarily including HTML, XML, and JSON which often derived from web-scraped data. This leaves domain-specific or less-studied markup languages, as well as proprietary formats in industry, underexplored. Addressing this gap requires the creation of specialized benchmarks for languages such as JATS, LaTeX, and Markdown to evaluate tasks like structure-aware question answering and format conversion. Moreover, model architectures must evolve towards structure-aware tokenization and training paradigms that explicitly leverage the hierarchical organization of markup, rather than treating markup as plain text sequences.
Understanding Table with Complex Structures. Semi-structured tables often exhibit irregular relationships and intricate nested structures, which remain difficult for current LLMs to fully capture. While recent approaches attempt to encode table structures and leverage the semantic reasoning capacity of LLMs, these methods are usually tailored to specific table types (e.g., financial tables). It is quite challenging to handle diverse semi-structured tables with complex structures. To address this, researchers have explored several directions. One line of work focuses on Structure-Aware Encodings, where the table’s hierarchical or graph-like organization (e.g., DOM trees for HTML tables or JSON schemas) is explicitly modeled and injected into LLMs. Another approach emphasizes Intermediate Representations, such as transforming semi-structured tables into canonical forms or domain-specific languages (DSLs) that better align with LLMs’ sequential reasoning ability. In addition, Retrieval-Augmented Methods leverage external knowledge bases or schema repositories to supplement missing semantics, enabling LLMs to ground their reasoning in richer contexts.
@@ 6.3 Unstructured Data
High-Level Chart Understanding. While significant progress has been made in basic chart recognition and data extraction, critical higher-order tasks such as chart-based fact checking, detection of visual or statistical misrepresentations (e.g., truncated axes, inconsistent scales, or fabricated trends), and cross-modal chart data verification remain largely unaddressed. These tasks demand not only accurate perception of visual elements but also reasoning about data semantics, statistical validity, and alignment with external knowledge or textual assertions. The absence of standardized benchmarks and evaluation protocols for such tasks further hinders systematic progress. We advocate for the development of comprehensive frameworks that support critical chart literacy—enabling systems to not just “read” charts, but to audit them. This includes creating datasets with annotated errors, factual inconsistencies, and deceptive visual practices, as well as designing models capable of joint visual-semantic reasoning to validate the integrity and truthfulness of the graph.
@@
Inadequate Multimodal Reasoning for Diverse Chart Types. Charts are inherently multimodal artifacts that fuse structured visual encodings (e.g., glyphs, spatial arrangements, color mappings) with natural language (titles, axis labels, captions). Current models often treat chart understanding as a vision-only or simplified vision-language problem, failing to capture the nuanced interplay between graphical conventions, domain-specific semantics, and contextual information. This limitation becomes especially pronounced when handling complex or non-standard chart types (e.g., small multiples, Sankey diagrams, or interactive dashboards). A holistic, end-to-end chart understanding pipeline that unifies low-level perception (e.g., element detection), mid-level parsing (e.g., data-value mapping), and high-level inference (e.g., trend interpretation or anomaly detection) is needed. Such a pipeline should be modular, chart-type-agnostic, and grounded in both visualization theory and real-world usage patterns, enabling robust performance across diverse downstream applications.
@@
Computational Efficiency of Video Analysis. Video is long-form sequential data; fixed sampling or full-frame processing consumes large compute and wastes capacity. Compressing frames/tokens often sacrifices fine-grained temporal cues, hurting localization and detail recognition. To address this, we can develop representations that preserve fine-grained visual detail while offering controllable context length (e.g., temporal separator tokens, dynamic sampling, learnable token merging, temporal positional encodings). Integrate temporal modeling and token-aggregation objectives into training to achieve a better trade-off between efficiency and fidelity.
@@
Video Temporal Localization with Semantic Coherence. Precise temporal localization requires detecting sparse, transient cues (e.g., action start/end, sudden events) and maintaining semantic consistency across long sequences. Current methods still struggle with event-boundary detection, timestamp reasoning, and preventing erroneous cross-time associations. To address this,we can improve fine-grained event detection and semantic continuity by incorporating temporally aware tokens and supervision objectives directly into LLM-friendly representations.
@@
Video Multimodal Fusion and Scalability. Effectively time-aligning and fusing visual, audio, and textual streams is both modeling and engineering challenging; high-quality, fine-grained human annotation is costly and hard to scale. Domain-specific tasks (emotion, gesture, anomaly detection, etc.) often rely on specialized labels and bespoke pipelines, which limits generalization. To address this,we can combine modular architectures (lightweight temporal adapters, instruction-aware selectors), multi-agent data curation/reviewer pipelines, and large-scale synthetic video–text generation to improve annotation efficiency, domain adaptation, and instruction-conditioned reasoning. Encourage task-centered cooperative agents that decompose complex video-understanding problems into simpler subtasks.
@@
Generalizing Multimodal Understanding for Documents. A key challenge beyond OCR is achieving deep multimodal understanding—integrating text, layout, and visuals—across diverse domains and languages. General-purpose models often fail on specialized document structures, such as dense scientific layouts or non-linear historical manuscripts, and struggle with language-specific conventions like right-to-left scripts. Addressing this requires document-native architectures (e.g., DocLLM) that emphasize layout over costly visual encoding, combined with layout-aware instruction tuning to teach multi-granularity structural reasoning, supported by large-scale, multilingual datasets for robust generalization.
@@
Efficiency in Large-Scale Document Processing. Processing large volumes of long-form documents, such as legal archives, presents a trade-off between computational efficiency and analytical depth. The limited context windows of most LLMs force chunking strategies that can sever long-range dependencies, while preserving fine-grained spatial details across thousands of pages remains a major technical hurdle. Future work should focus on adaptive context management and hierarchical document representations. Combining these with progressive reasoning strategies, where a model first builds a coarse understanding before focusing on details, can create a better balance between computational cost and high-fidelity analytical accuracy.
@@
Accuracy and Robustness in Program Vulnerability Detection. While LLMs show promise in software vulnerability detection, the accuracy of state-of-the-art methods remains unsatisfactory, with reported success rates around 67.6% for detection and only 20% for automated repair. The number of vulnerabilities continues to grow, with over 25,000 new CVEs reported in 2024 alone. Moreover, the performance of deep learning-based methods has been shown to degrade significantly on rigorously validated datasets and in the face of simple code modifications. The development of novel frameworks that integrate Code Property Graphs (CPGs) with LLMs presents a promising path. Techniques like CPG-guided code slicing can reduce code size while preserving vulnerability-relevant context, leading to significant improvements in F1 scores for inter-procedural vulnerability detection.
@@
Program Analysis Complexity and Scope Limitations. The inherent complexity of vulnerabilities poses a significant challenge, with inter-procedural flaws being substantially harder to detect than intra-procedural ones. LLMs often struggle with uncommon CWE types, and their performance degrades when vulnerabilities span multiple code units. Most current LLM-based solutions are limited to the function level, a narrow scope that questions their effectiveness for class-level or repository-level analysis. Future research should focus on developing LLM architectures that support multi-level code analysis, scaling from the function level to project and enterprise levels. Integrating Retrieval-Augmented Generation (RAG) techniques is crucial for effectively handling the vast contextual information present in large-scale codebases.
@@
Reliability in Code Generation and Repair. Although LLMs offer advantages in generality and interpretability—providing intermediate reasoning for their findings—they still face significant challenges in terms of trust and accuracy, particularly for automated code repair. Establishing mechanisms for trust and collaboration with developers is essential. LLMs should be designed to signal uncertainty rather than providing potentially inaccurate outputs. Furthermore, developing multi-agent collaborative frameworks, with specialized agents for code analysis, vulnerability detection, and repair, could lead to more robust and reliable systems.
@@
High Quality 3D Modality Representation. 3D data, including point clouds, meshes, voxels, implicit fields, and NeRFs, vary in sampling, topology, and attributes. Converting them into LLM-ready inputs requires trade-offs, where aggressive simplification loses geometric detail, while token explosion exceeds context limits, leading to inconsistent cues and poor generalization. This can be mitigated with compact, trainable encoders that produce a small set of informative tokens, such as vector-quantized mesh tokens, learned point-feature aggregators, and geometry-aware positional encodings. Jointly training these encoders with lightweight adapters using multiview consistency, reprojection, and geometric-accuracy losses preserves topology and metric cues while staying within LLM context budgets.
@@
Accurate 3D Geometric Grounding. LLMs lack native metric reasoning and struggle with coordinate frames, units, and precise spatial numerics. The scarcity of richly labeled 3D datasets limits supervised learning for distances, poses, and geometric constraints, so systems often produce plausible qualitative descriptions but fail on quantitative queries, affecting navigation, manipulation, simulation, and safety-critical tasks. This can be addressed using domain adapters that encode coordinate conventions, units, and regression heads for distances/poses, trained with metric supervision from simulators/renderers or differentiable-rendering losses, and evaluated on numerically grounded benchmarks to ensure quantitative fidelity.
@@
Pluggable Modality Extensibility. A key challenge for heterogeneous data agents is integrating new data types without redesigning or retraining the system. Traditional architectures are rigid, making it difficult to handle emerging modalities like audio, video, or sensor data. Future work should focus on modular and extensible frameworks with independent encoders or adapters, standardized integration interfaces, adaptive weighting for multiple modalities, and self-supervised alignment to ensure compatibility with existing representations. Such approaches will enhance scalability, maintainability, and robustness in increasingly complex and evolving data.
High-Level Heterogeneous Data Analysis. A key challenge for heterogeneous data agents is performing high-level analysis, encompassing reasoning, summarization, and cross-modal inference beyond simple retrieval. This requires understanding interrelationships among modalities, capturing context, and generating actionable insights. Existing systems, such as XMODE nooralahzadeh2024explainable and wang2024interactive , offer partial solutions via step-wise query decomposition or adaptive modality weighting but still struggle with large-scale or disorganized datasets, limiting fully automated, context-aware analysis. Future research should develop agents capable of robust, end-to-end reasoning across heterogeneous modalities. Promising directions include using multimodal embeddings for joint inference, integrating symbolic and neural reasoning for interpretability, and employing adaptive task planning to decompose complex queries and synthesize results across modalities for comprehensive real-world data analysis.